# Surprisingly High Redundancy in Electronic Structure Data

### v1

@@ IV.1 Machine learning map for electron density prediction
ML approaches for predicting electron densities use descriptors that encode the atomic configuration of a local spatial neighborhood as input. Effective descriptors include interatomic distances and angles, as well as the Superposition of Atomic Densities (SAD), both of which have been successful in capturing local atomic environments [20, 27, 28, 29]. The output of these ML models, the electron density, is commonly represented in one of two ways: (1) as sums of atom-centered basis functions [90, 91], which offer computational efficiency but may struggle with complex densities, and (2) as grid-based predictions [19, 21], which provide higher accuracy but require significant computational resources due to fine-mesh evaluations.
@@
As proposed in our recent contribution  [28], we adopt the position vectors of atoms in body-attached local frames and the atomic species as the descriptors. This formulation ensures that the descriptors are invariant with respect to rotations, translations and species permutations of the system. Moreover, as shown in  [28] the descriptors maintain approximately the same descriptor-vector size even if the number of species in the system increases, thus ensuring efficient representation of atomic neighborhoods across chemical space. The local frame of reference for calculating the descriptors is obtained using Principal Component Analysis (PCA) of an atomic neighborhood consisting of $M$ atoms. The components of the position vector ($\textbf{r}_{j}$) of the $j$-th atom in the local reference frame are denoted by $(x,y,z)$. The species information is encoded by using the atomic number of the species. The atomic number of the $j$-th atom is denoted as $Z_{j}$. Thus the descriptors for $i$-th grid point ($\mathcal{D}_{i}\in\mathbb{R}^{5M}$) are given as,
$\mathcal{(}\mathcal{D}_{i})^{j}=\bigg{\{}Z_{j},\|\textbf{r}_{j}\|,\frac{x}{\|\textbf{r}_{j}\|},\frac{y}{\|\textbf{r}_{j}\|},\frac{z}{\|\textbf{r}_{j}\|}\bigg{\}};\;j=1,\cdots,M$ (1)
@@ IV.2 Pruning strategies
We briefly recapitulate the pruning strategies used in this work in addition to the random pruning: first, the GraNd score-based pruning [65, 62], followed by a Coverage-centric Coreset Selection (CCS) method [68] applied to the GraNd scores.
@@ IV.2.2 GraNd score based pruning:
We consider a supervised regression setting, where the training set is $S=(x_{i},y_{i})_{i=1}^{N}$, drawn i.i.d. from an underlying distribution $P$. Here, $x_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\mathbb{R}^{+}$ denote the descriptors at $i$-th grid point and the electronic charge density at that grid point, respectively. Furthermore, $d$ is the dimension of the atomic-neighborhood descriptors. For a fixed neural network, let the w be the weights and $l$ be the loss of the neural network at a given step. Let $\textbf{w}_{0}$, $\textbf{w}_{1}$, … , $\textbf{w}_{T}$ be iteratively updated weights of the neural network, where, for some sequence of minibatches $S_{0}$, $S_{1}$, …, $S_{T-1}$ $\subseteq S$ of size $M$, we have $\textbf{w}_{t}=\textbf{w}_{t-1}-\eta\sum_{(x,y)\epsilon S_{t-1}}g_{t-1}(x,y)$. Here $g_{t}(x,y)=\nabla_{\textbf{w}_{t}}l(x,y)$, $t=1,...,T$, and $\eta$ is the learning rate. The iterations can be perceived as time evolution, with the time derivative of the loss denoted by, $\Delta_{t}((x,y),S_{t})=-\frac{dl}{dt}$. Due to the random initialization of the the neural network, the weight vector at time $t>0$, $\mathbf{w}_{t}$, is considered as a random variable.
GraNd score: The GraNd score of a training example $(x,y)$ is defined at time $t$ as $\chi_{t}(x,y)=\mathbb{E}_{\mathbf{w}_{t}}||g_{t}(x,y)||_{2}\,$, [65]. It represents the impact of removal of an example $\left(x_{j},y_{j}\right)$ from the training set ($S$) on the change in the loss for other examples. The change in the time derivative of the loss of any example $(x^{*},y^{*})$ resulting from the removal of a training example $(x_{j},y_{j})$ from the dataset $S$ is bounded by the GraNd score associated with the removed example. Formally, for all $(x^{*},y^{*})$, there exists a constant $c$ such that $\big{\|}\Delta_{t}((x^{*},y^{*});S)-\Delta_{t}((x^{*},y^{*});S_{-j})\big{\|}\leq c\,\|g_{t}(x_{j},y_{j})\|$. Where $S_{-j}=S\setminus\{(x_{j},y_{j})\}$ and $\Delta_{t}((x^{*},y^{*});S)=-\frac{d\ell}{dt}$. The constant $c$ does not depend on the training examples, $(x,y)$s. Therefore, the contribution of an example $(x,y)$ to change in the loss of any other example $(x^{*},y^{*})$ is bounded by the gradient norm, $\|g_{t}(x_{j},y_{j})\|$, at any time $t$, given the weights $\mathbf{w}_{t}$. In practice, the gradient of the loss, $l$, is computed with respect to the model’s weights $\mathbf{w}_{t}$ at epoch $t$ for each training example. The GraNd score is obtained by computing the expectation over several training runs of $L_{2}$ norm of the gradient. The examples are ranked based on their GraNd scores, with higher scores indicating greater difficulty in learning.
@@ IV.2.3 Coverage-centric Coreset Selection
The purpose of coreset selection is to choose a subset ($\mathcal{S^{\prime}}$) of the data ($\mathcal{S}$) that minimizes the loss on the test set [68]. This selection process can be represented through the following optimization problem [69]:
@@
where $l$ is the loss function, $h_{\mathcal{S}^{\prime}}$ is the model trained with the coreset $\mathcal{S}^{\prime}$, and $f$ is the pruning factor. In other words, $\mathcal{S}^{\prime}$ is a subset of $\mathcal{S}$ which has a size at most $(1-f)$ of $S$, and which minimizes the expected test loss over the distribution ${P}$.
In a classical coverage setting, a set $\mathcal{S^{\prime}}$ is called an r-cover set of $\mathcal{S}$ if a ball of radius $r$ centered at each data point of $\mathcal{S^{\prime}}$ covers the entire $\mathcal{S}$. Zheng et al. [68] extended the classical geometric set cover to cover a probability distribution—referred to as the density-based distribution cover. The associated coreset selection method is called the Coverage-centric Coreset Selection (CCS). They also introduced the concept of a partial cover percentage $p$, of the probability distribution $P$, instead of the full cover.
@@
[…]
Here $B_{d}(\textbf{x},r)=\{\textbf{x}^{\prime}\in X:d(\textbf{x},\textbf{x}^{\prime})\leq r\}$ denotes a $r$-radius ball centered at x, $p$ quantifies coverage (in percentage), and $\mu$ represents the probability measure associated with the probability distribution $P$. The probability density information over the input space is reflected in the measure $\mu$. The percentage of coverage $p$ increases with the increasing radius $r$ for a given set $\mathcal{S}$ and a distribution $P$. Given a coverage percentage $p$ and a set $\mathcal{S}$, the minimum covering radius $r$ is the smallest value such that $\mathcal{S}$ achieves at least $p$ percentage coverage under distribution $P$. In practice, the distribution $P$ is unknown and so an approximate estimation of the coverage is required. Thus, the area under the p-r curve ($\text{AUC}_{pr}$), is proposed as a metric to assess the coverage of a coreset [68]. The quantity $\text{AUC}_{pr}$ defined as the expectation of the minimum distance between the examples that follows the distribution $P$ and the data in $\mathcal{S}$: $\text{AUC}_{pr}(\mathcal{S})=\mathbb{E}_{x\sim P}\left[\min_{x^{\prime}\in\mathcal{S}}d(\textbf{x}^{\prime},\textbf{x})\right]$. The lower the value of the $\text{AUC}_{pr}$, the better the coverage by the coreset.
@@
CCS steps: Consider a dataset $\mathcal{S}$ = $\{(x_{i},y_{i},s_{i})\}_{i=1}^{n}$ where $s_{i}$ is the importance score for i-th example $(x_{i},y_{i})$. First, CCS selects a pruning rate $\beta$, a tunable hyperparameter representing the percentage of hard examples to remove. Pruning $\mathcal{\beta}\,\%$ of hard examples from $\mathcal{S}$ yields $\mathcal{S}_{o}$. Second, CCS partitions $\mathcal{S}_{o}$ into $k$ strata based on importance (GraNd) scores, using fixed-width score intervals. While each stratum has the same score range, the number of examples per stratum may vary. Third, given the pruning factor $f$, the desired coreset size is determined as $m=(1-f)\,n$. This total budget $m$ is distributed evenly across the $k$ strata (another hyperparameter), allocating $m/k$ samples per stratum. If a stratum has fewer examples than its budget, all its examples are included. The final coreset ($\mathcal{S^{\prime}}$) is formed by combining the sampled examples from all strata.
@@
SH, PT and SP worked on developing the framework for pruning and other machine learning aspects. ST and AG worked on the DFT data generation and post-processing calculations. ASB and SG were involved in conceptualization, methodological design, supervision, and securing funding/resources. All authors contributed to writing the manuscript.
@@
Codes supporting the findings of this study are available from the corresponding author upon reasonable request.
@@ Acknowledgements.
This work was primarily supported by grant DE-SC0023432 funded by the U.S. Department of Energy, Office of Science. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231, using NERSC awards BES-ERCAP0033206, BES-ERCAP0025205, BES-ERCAP0025168, and BES-ERCAP0028072. SG acknowledges the grant 2442313 funded by the U.S. National Science Foundation. The authors would like to thank UCLA’s Institute for Digital Research and Education (IDRE), the MRI GPU cluster at MTU for making available some of the computing resources used in this work. The authors acknowledge the use of the GPT-4o (OpenAI) model to polish the language and edit grammatical errors in some sections of this manuscript. The authors subsequently inspected, validated and edited the text generated by the AI model, before incorporation.

### v2

@@ IV.1 Machine learning map for electron density prediction
ML approaches for predicting electron densities use descriptors that encode the atomic configuration of a local spatial neighborhood as input. Effective descriptors include interatomic distances and angles, as well as the Superposition of Atomic Densities (SAD), both of which have been successful in capturing local atomic environments [chandrasekaran2019solving, pathrudkar2024electronic, pathrudkar2024HEA, li2025image]. The output of these ML models, the electron density, is commonly represented in one of two ways: (1) as sums of atom-centered basis functions [grisafi2018transferable, fabrizio2019electron], which offer computational efficiency but may struggle with complex densities, and (2) as grid-based predictions [zepeda2021deep, fiedler2023predicting], which provide higher accuracy but require significant computational resources due to fine-mesh evaluations.
@@
As proposed in our recent contribution  [pathrudkar2024HEA], we adopt the position vectors of atoms in body-attached local frames and the atomic species as the descriptors. This formulation ensures that the descriptors are invariant with respect to rotations, translations and species permutations of the system. Moreover, as shown in  [pathrudkar2024HEA] the descriptors maintain approximately the same descriptor-vector size even if the number of species in the system increases, thus ensuring efficient representation of atomic neighborhoods across chemical space. The local frame of reference for calculating the descriptors is obtained using Principal Component Analysis (PCA) of an atomic neighborhood consisting of $M$ atoms. The components of the position vector ($\textbf{r}_{j}$) of the $j$-th atom in the local reference frame are denoted by $(x,y,z)$. The species information is encoded by using the atomic number of the species. The atomic number of the $j$-th atom is denoted as $Z_{j}$. Thus the descriptors for $i$-th grid point ($\mathcal{D}_{i}\in\mathbb{R}^{5M}$) are given as,
$\mathcal{D}_{i}=\bigg\{Z_{j},\|\textbf{r}_{j}\|,\frac{x}{\|\textbf{r}_{j}\|},\frac{y}{\|\textbf{r}_{j}\|},\frac{z}{\|\textbf{r}_{j}\|}\bigg\}_{j=1,\cdots,M}$ (1)
@@ IV.2 Pruning strategies
We briefly recapitulate the pruning strategies used in this work in addition to the random pruning: first, the GraNd score-based pruning [paul2021deep, sorscher2022beyond], followed by a Coverage-centric Coreset Selection (CCS) method [zheng2023_CCS] applied to the GraNd scores.
@@ IV.2.2 GraNd score based pruning:
We consider a supervised regression setting, where the training set is $S=(x_{i},y_{i})_{i=1}^{N}$, drawn i.i.d. from an underlying distribution $P$. Here, $x_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\mathbb{R}^{+}$ denote the descriptors at $i$-th grid point and the electronic charge density at that grid point, respectively. Furthermore, $d$ is the dimension of the atomic-neighborhood descriptors. For a fixed neural network, let the w be the weights and $l$ be the loss of the neural network at a given step. Let $\textbf{w}_{0}$, $\textbf{w}_{1}$, … , $\textbf{w}_{T}$ be iteratively updated weights of the neural network, where, for some sequence of minibatches $S_{0}$, $S_{1}$, …, $_{T-1} S\subseteq S$ of size $M$, we have $\textbf{w}_{t}=\textbf{w}_{t-1}-\eta\sum_{(x,y)\epsilon S_{t-1}}g_{t-1}(x,y)$. Here $g_{t}(x,y)=\nabla_{\textbf{w}_{t}}l(x,y)$, t=1,,Tt=1,...,T, and $\eta$ is the learning rate. The iterations can be perceived as time evolution, with the time derivative of the loss denoted by, $\Delta_{t}((x,y),S_{t})=-\frac{dl}{dt}$. Due to the random initialization of the the neural network, the weight vector at time $t>0$, $\mathbf{w}_{t}$, is considered as a random variable.
[…]
The change in the time derivative of the loss of any example $(x^{*},y^{*})$ resulting from the removal of a training example $(x_{j},y_{j})$ from the dataset $S$ is bounded by the GraNd score associated with the removed example. Formally, for all $(x^{*},y^{*})$, there exists a constant $c$ such that $\big\|\Delta_{t}((x^{*},y^{*});S)-\Delta_{t}((x^{*},y^{*});S_{-j})\big\|\leq c\,\|g_{t}(x_{j},y_{j})\|$. Where $S_{-j}=S\setminus\{(x_{j},y_{j})\}$ and $\Delta_{t}((x^{*},y^{*});S)=-\frac{d\ell}{dt}$. The constant $c$ does not depend on the training examples, $(x,y)$s. Therefore, the contribution of an example $(x,y)$ to change in the loss of any other example $(x^{*},y^{*})$ is bounded by the gradient norm, $\|g_{t}(x_{j},y_{j})\|$, at any time $t$, given the weights $\mathbf{w}_{t}$. In practice, the gradient of the loss, $l$, is computed with respect to the model’s weights $\mathbf{w}_{t}$ at epoch $t$ for each training example. The GraNd score is obtained by computing the expectation over several training runs of $L_{2}$ norm of the gradient. The examples are ranked based on their GraNd scores, with higher scores indicating greater difficulty in learning.
@@ IV.2.3 Coverage-centric Coreset Selection
The purpose of coreset selection is to choose a subset ($\mathcal{S^{\prime}}$) of the data ($\mathcal{S}$) that minimizes the loss on the test set [zheng2023_CCS]. This selection process can be represented through the following optimization problem [sener2018coreset]:
@@
where $l$ is the loss function, $h_{\mathcal{S}^{\prime}}$ is the model trained with the coreset $\mathcal{S}^{\prime}$, and $f$ is the pruning factor. In other words, $\mathcal{S}^{\prime}$ is a subset of $\mathcal{S}$ which has a size at most (1f)(1-f) of $S$, and which minimizes the expected test loss over the distribution P{P}.
In a classical coverage setting, a set $\mathcal{S^{\prime}}$ is called an r-cover set of $\mathcal{S}$ if a ball of radius $r$ centered at each data point of $\mathcal{S^{\prime}}$ covers the entire $\mathcal{S}$. Zheng et al. [zheng2023_CCS] extended the classical geometric set cover to cover a probability distribution—referred to as the density-based distribution cover. The associated coreset selection method is called the Coverage-centric Coreset Selection (CCS). They also introduced the concept of a partial cover percentage $p$, of the probability distribution $P$, instead of the full cover.
@@
$\int_{X}\mathbf{1}_{\cup_{x\in\mathcal{S}}B_{d}(\textbf{x},r)}(\textbf{x})\,d\mu(\textbf{x})=p$
Here $B_{d}(\textbf{x},r)=\{\textbf{x}^{\prime}\in X:d(\textbf{x},\textbf{x}^{\prime})\leq r\}$ denotes a $r$-radius ball centered at x, $p$ quantifies coverage (in percentage), and $\mu$ represents the probability measure associated with the probability distribution $P$. The probability density information over the input space is reflected in the measure $\mu$. The percentage of coverage $p$ increases with the increasing radius $r$ for a given set $\mathcal{S}$ and a distribution $P$. Given a coverage percentage $p$ and a set $\mathcal{S}$, the minimum covering radius $r$ is the smallest value such that $\mathcal{S}$ achieves at least $p$ percentage coverage under distribution $P$. In practice, the distribution $P$ is unknown and so an approximate estimation of the coverage is required. Thus, the area under the p-r curve ($\text{AUC}_{pr}$), is proposed as a metric to assess the coverage of a coreset [zheng2023_CCS]. The quantity $\text{AUC}_{pr}$ defined as the expectation of the minimum distance between the examples that follows the distribution $P$ and the data in $\mathcal{S}$: $\text{AUC}_{pr}(\mathcal{S})=\mathbb{E}_{x\sim P}\left[\min_{x^{\prime}\in\mathcal{S}}d(\textbf{x}^{\prime},\textbf{x})\right]$. The lower the value of the $\text{AUC}_{pr}$, the better the coverage by the coreset.
@@
CCS steps: Consider a dataset $\mathcal{S}$ = $\{(x_{i},y_{i},s_{i})\}_{i=1}^{n}$ where $s_{i}$ is the importance score for i-th example $(x_{i},y_{i})$. First, CCS selects a pruning rate $\beta$, a tunable hyperparameter representing the percentage of hard examples to remove. Pruning $\mathcal{\beta}\,\%$ of hard examples from $\mathcal{S}$ yields $\mathcal{S}_{o}$. Second, CCS partitions $\mathcal{S}_{o}$ into $k$ strata based on importance (GraNd) scores, using fixed-width score intervals. While each stratum has the same score range, the number of examples per stratum may vary. Third, given the pruning factor $f$, the desired coreset size is determined as m=(1f)nm=(1-f)\,n. This total budget $m$ is distributed evenly across the $k$ strata (another hyperparameter), allocating $m/k$ samples per stratum. If a stratum has fewer examples than its budget, all its examples are included. The final coreset ($\mathcal{S^{\prime}}$) is formed by combining the sampled examples from all strata.
@@
SH, PT and SP worked on developing the framework for pruning and other machine learning aspects. ST and AG worked on the KS-DFT data generation and post-processing calculations. ASB and SG were involved in conceptualization, methodological design, supervision, and securing funding/resources. All authors contributed to writing the manuscript.
@@
Codes supporting the findings of this study are available from the corresponding authors upon reasonable request.
Materials & Correspondence
Correspondence and requests for materials should be addressed to:
Amartya S. Banerjee (email: asbanerjee@ucla.edu), and
Susanta Ghosh (email: susantag@mtu.edu)
Funding
This work was primarily supported by grant DE-SC0023432 funded by the U.S. Department of Energy, Office of Science. This work was partially supported by the National Science Foundation under Grant No. 2442313 and through the UC National Laboratory Fees Research Program of the University of California, Grant Number L25CR9003.
@@ Acknowledgements.
This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231, using NERSC awards BES-ERCAP0033206, BES-ERCAP0025205, BES-ERCAP0025168, and BES-ERCAP0028072. The authors would like to thank UCLA’s Institute for Digital Research and Education (IDRE), the MRI GPU cluster at MTU for making available some of the computing resources used in this work. The authors acknowledge the use of the GPT-5 (OpenAI) model to polish the language and edit grammatical errors in some sections of this manuscript. The authors subsequently inspected, validated and edited the text generated by the AI model, before incorporation.
V Supplemental Materials
V.1 Error Metrics
The normalized root mean square error (NRMSE) of the ML predicted electron density is defined as
$NRMSE=\frac{\sqrt{\frac{1}{N}\sum_{i=1}^{N}(\rho^{i}_{DFT}-\rho^{i}_{ML})^{2}})}{\rho_{DFT}^{max}-\rho_{DFT}^{min}}$ (3)
where N is the number of grid-points for each snapshot. At $i$-th grid point, $\rho^{i}_{DFT}$ represents the electron density calculated using KS-DFT while $\rho^{i}_{ML}$ represents the electron density predicted by the ML models. $\rho_{DFT}^{max}$ and $\rho_{DFT}^{min}$ are the maximum and minimum values of electron density across all grid-points of that snapshot.
Let $u$ is the error field of the ML-predicted electron density, given by $u(x)=\rho_{DFT}(x)-\rho_{ML}(x)$, with $x\in\Omega$. The $H^{1}$ seminorm of the scalar error field $u(x)$ over a domain $\Omega$ is defined as $|u|_{H^{1}(\Omega)}=\left(\int_{\Omega}|\nabla u(x)|^{2}\,dx\right)^{1/2}$. The $H^{1}$ norm is defined as $\|u\|_{H^{1}(\Omega)}=\left(\|u\|_{L^{2}(\Omega)}^{2}+|u|_{H^{1}(\Omega)}^{2}\right)^{1/2}$. These quantities are well defined whenever, $u$ lies in the Sobolev space $H^{1}(\Omega)$.
In the paper, the energy prediction errors are defined as the absolute differences (in Hartree per atom) between KS-DFT energies and the energies obtained by postprocessing the ML-predicted electron density.
V.2 Postprocessing of ML predicted electron density
For application purposes, it is of interest to compute downstream properties from the ML-predicted electron density. Since the Kohn-Sham total ground-state energy is a key property that can be computed from the electron density, it can serve as a useful test for the downstream accuracy of the ML model. We postprocess the density predictions to obtain the corresponding total ground-state energies as follows. The predicted electron density is first rescaled by the total number of electrons; this serves to ensure that the total electronic charge is consistent [pathrudkar2024electronic, briling2021impact, alred2018machine, pathrudkar2022machine]. The scaled electron density $\rho^{\text{scaled}}$ at grid point r is obtained from the predicted electron density $\rho^{\text{ML}}$ via:
$\rho^{\text{scaled}}\left(\textbf{r}\right)=\rho^{\text{ML}}(\textbf{r})\frac{N_{\text{e}}}{\displaystyle\int_{\Omega}\rho^{\text{ML}}(\textbf{r})d\textbf{r}}\,.$ (4)
Here $\Omega$ is the periodic supercell used in the calculations, and $N_{\text{e}}$ is the number of electrons in the system. The rescaled electron density $\rho^{\text{scaled}}$ is then used to set up the Kohn-Sham Hamiltonian and a single diagonalization step is performed by using the same electronic structure calculation framework, as was used in the data generation process [xu2021sparc, xu2020m, ghosh2017sparc]. Thereafter, the energy of the system is computed using the Harris-Foulkes formula [harris1985simplified, foulkes1989tight]. The difference between the energy obtained from the ML-predicted electron density and that obtained from the original DFT density serve as the energy error reported in this work.
V.3 Experimental Details
Dataset details: We use sufficiently large datasets to achieve high accuracy in ML for all four material systems. The KS-DFT dataset for the aluminum system consists of 200 training snapshots (216,000 grid points each), totaling $43.2\times 10^{6}$ grid points, obtained from a 32-atom simulation cell; the test set includes 60 snapshots from six temperatures. The dataset for water includes 112 training snapshots (456,533 grid points each, totaling $51.1\times 10^{6}$ grid points), and 193 test snapshots from four temperatures, obtained from a 32-atom simulation cell. The dataset for SiGeSn system, taken from 45 compositions, consists of 216 training snapshots (totaling $38.6\times 10^{6}$ grid points) and test set includes 90 test snapshots from 45 compositions, obtained from a 64-atom simulation cell. The dataset for CrFeCoNi system, taken from 69 compositions, consists of 140 training snapshots (totaling $40.9\times 10^{6}$ grid points) and test set includes 207 test snapshots from four different temperatures and 69 compositions, also from a 32-atom simulation cell.
Neural Network: We employ a deep neural network with 12 hidden layers, each containing 900 nodes, using GELU activation [hendrycks2016gaussian] in the hidden layers and a ReLU activation in the output layer. A learning rate of 0.001 is used across all cases, with the AdamW optimizer [kingma2014adam, loshchilov2017decoupled].
GraNd score computation: The epoch 10, 30, 10 and 10 are chosen for GraNd score computation for Aluminum, water, SiGeSn and CrFeCoNi respectively. Figure 8 shows the errors on the test dataset for Aluminum, water, SiGeSn, and CrFeCoNi for pruned subsets, for scores computed at different times in training. The Figure 8 shows that the computing scores early in training is sufficient to identify important examples, aligning with the findings of [paul2021deep].
Hyper-parameters for CCS: There are two hyper-parameters for the CCS method. a) the hard cut-off rate ($\beta$) and b) the number of strata ($k$). The optimal $\beta$ generally increases with the pruning rate. The performance of various CCS-based coreset constructed from varying $\beta$ is shown in Figure 10. The error in charge density prediction for various number of strata, $k$ is shown in Figure 9. The error is not very sensitive to the that the number of strata($k$), which is consistent with the findings of Zheng et al. [zheng2023_CCS] for image classification datasets. Hence, in this work, the number of strata($k$) being used is 100 all four material systems for all pruning factors.
Figure 8: Choice of score computation epoch: The error in electron density prediction for (a) Aluminum(20% pruning), (b) Water(40% pruning) (c) SiGeSn(20% pruning) and (d) CrFeCoNi(40% pruning) after computing the GraNd score at different times in training.
Figure 9: Choice of number of strata($k$):The error in electron density prediction with various numbers of strata for Aluminum system for various pruning factors for CCS method.
Figure 10: Choice of hard cut-off rate($\beta$): The error in electron density prediction with varying $\beta$ for (a) Aluminum, (b) water (c) SiGeSn (d) CrFeCoNi for the CCS based pruning. It is observed that the obtained NRMSE with pruned dataset vary with the choice of $\beta$. For higher pruning rate, a higher $\beta$ may be optimal. The points with $\times$ markers failed to beat random pruning.
V.4 Quantifying the coverage of a specific distribution by a subset
The GraNd score and other state-of-the-art importance-based pruning methods often perform poorly at higher pruning rates because of poor data coverage.
We have evaluated the $\text{AUC}_{pr}$ by computing the expected minimum distance from the data in the validation dataset to the data in coreset.
Aluminum
Pruning $\text{AUC}_{pr}$ NRMSE
Rate GraNd CCS GraNd CCS
90% 8.396 8.30 $58\times 10^{-2} 1.1$$15\times 10^{-2}$
99% 8.727 8.584 $2.55\times 10^{-2}$ $1.3\times 10^{-2}$
Water
Pruning $\text{AUC}_{pr}$ NRMSE
Rate GraNd CCS GraNd CCS
90% 23.18 22.93 $3.1\times 10^{-2}$ $1.09\times 10^{-3}$
99% 25.29 24.89 $1.11\times 10^{-1}$ $1.9\times 10^{-3}$
SiGeSn
Pruning $\text{AUC}_{pr}$ NRMSE
Rate GraNd CCS GraNd CCS
90% 76.57 76.30 $2.55\times 10^{-2}$ $8.27\times 10^{-3}$
99% 81.91 80.537 $8.27\times 10^{-2}$ $1.17\times 10^{-2}$
CrFeCoNi
Pruning $\text{AUC}_{pr}$ NRMSE
Rate GraNd CCS GraNd CCS
90% 13.074 13.039 $4.4\times 10^{-2}$ $2.4\times 10^{-3}$
99% 13.416 13.335 $1.27\times 10^{-1}$ $2.88\times 10^{-3}$
Table 1: Comparison of $\text{AUC}_{pr}$ and NRMSE between GraNd and CCS method for different pruning percentages
From the Table 1, we can see that, for 90% and 99% pruning percentage, the CCS has lower $\text{AUC}_{pr}$ than the GraNd method. It confirms that the CCS method based coreset indeed selects a subset that ensures a better data coverage, as evidenced by its lower errors (NRMSE).
V.5 Generalizability
In this section, we provide additional results that demonstrate the generalizability of ML models trained on pruned datasets. The error in electron density prediction for defects in the 32-atoms aluminum system is shown in Figure 11. The generalization error of the GraNd method increases rapidly beyond 8080% pruning, whereas random pruning maintains high accuracy up to 9090% pruning. The CCS method supports a remarkably high pruning factor (9999%) without any significant loss in generalization performance
The errors in electron density prediction for handcrafted checkerboard SiGeSn alloy systems is shown in Figure 12. Checkerboard systems of 216 atoms and 64 atoms were considered. The generalization performance for GraNd worsen rapidly beyond 6060%, whereas random pruning maintains high accuracy up to 8080% pruning factor. However, CCS maintains near-original performance until 9999% pruning.
For CrFeCoNi system with defects, the errors in electron density prediction for defects is shown in Figure 13. The generalization performance for GraNd worsen rapidly beyond 4040%, whereas random pruning maintains high accuracy up to 8080% pruning factors. However, CCS maintains near-original performance until 9999% pruning.
Figure 11: Generalizability tests for Aluminum system: The error in the electron density prediction for various pruning factors for monovacancies for 32-atoms Aluminum system. The shaded region represents the range (maximum to minimum) of three ML models around the mean, shown by a solid line.
Figure 12: Generalizability tests for SiGeSn system: The error in the electron density prediction for various pruning factors for SiGeSn handcrafted checkerboard systems. The shaded region represents the range (maximum to minimum) of three ML models around the mean, shown by a solid line.
Figure 13: Generalizability tests for CrFeCoNi system: The error in the electron density prediction for various pruning factors for (a) mono- (b) di-vacancy defects for 32-atoms CrFeConi system. The shaded region represents the range (maximum to minimum) of three ML models around the mean, shown by a solid line. Figures(c) and (d) shows the magnified view of the Figures(a) and (b)
Figure 14: Error in energy for generalizability tests for CrFeCoNi Monovacancy: The error in the energy from the ML-predicted electron density for the original dataset, 90% and 99% CCS-based dataset.
Figure 15: Model generalizability to unseen CrFeCoNi monovacancy defect configurations: Two-dimensional(2D) slices of electron density obtained by KS-DFT, ML model trained on the original dataset, 90% CCS and 99% CCS based pruned dataset. The electron density along the solid line is also compared with DFT and it shows remarkable agreement with the DFT. The unit of electron density is $\text{e}\cdot\text{Bohr}^{-3}$, where e denotes the electronic charge.
We also compute the total energy of the system from the ML-predicted electron density, via postprocessing, for these generalizability tests. The total energy errors (Hartree/atom) relative to the KS-DFT calculations, are shown in Figure 14, which shows that the even after 99% pruning through CCS, the ML predictions are at the chemically accuracy levels. It demonstrates that the CCS-based pruning, even at two orders of magnitude reduction, retains a remarkably high level of the original model’s generalization capability.
V.6 Timing Comparison
Figure 16: Training time comparison between the ML model trained on the original dataset, 90% CCS and 99% CCS based pruned dataset
Pruned subsets speed up the training process significantly, as shown in Figure 16. Training on $90\%$ CCS-pruned subsets reduces training time (compared to the original dataset) by $70\%$, $84\%$, $74\%$, and $50\%$ for Aluminum, water, SiGeSn, and CrFeCoNi systems respectively, while matching the original models’ predictive performance. $99\%$ CCS-based pruned subsets reduces the training time further, while also suffering nominal increase in NRMSE.
V.7 Additional Results
A complete and unscaled view of the error versus pruning factor for water, SiGeSn, and CrFeCoNi, (corresponding to Figures 5 and 6 of the main text) is provided in Figure 17, Figure 18, and Figure 19 for completeness. It shows that the GraNd method becomes too erroneous beyond 60% pruning for water, SiGeSn and CrFeCoNi. However, the random pruning provides high accuracy up to 90%.
Figure 18: SiGeSn system: (a) Error in the predicted electron density for various pruning factors. (b) Error in energy obtained from the predicted electron density for various pruning factors.
Figure 19: CrFeCoNi system: (a) Error in the predicted electron density for various pruning factors. (b) Error in energy obtained from the predicted electron density for various pruning factors.
V.7.1 Isosurface Plots
The 3D electron density fields are shown as iso-surface plots for the four materials in Figures 20, 22, 24, and 26. The overall pattern of electron density iso-surfaces is well captured by the original dataset, 90%, and 99% CCS based pruned datasets.
The corresponding 3D error fields are shown in Figures 21, 23, 25, and 27. Although the error iso-surfaces closely resemble those of the original datasets even after 9090% and 9999% pruning, a slight yet progressive increase in error is observed as pruning levels increase.
Figure 20: Aluminum system at 1500 K. Iso-surfaces of electron density obtained by KS-DFT (a) and ML model for original (unpruned) (b), 9090% CCS (c), and 9999% CCS (d).
Figure 21: Aluminum system at 1500 K. Iso-surfaces of errors in electron density obtained by ML model for original (a), 90% CCS (b), and 99% CCS(c) respectively.
Figure 22: Water system at 600K: Iso-surfaces of electron density (at 0.1 e) obtained by DFT (a) and ML model for original (b), 90% CCS (c), and 99.5% CCS (d). The red and pink spheres represent the Oxygen and Hydrogen atoms.
Figure 23: Water system at 600K: Iso-surfaces of errors in electron density (at 0.0075 e) obtained by ML model for original (unpruned) (a), 90% CCS (b), and 99.5% CCS (c) respectively.
Figure 24: $\text{Si}_{50}\text{Ge}_{12.5}\text{Sn}_{37.5}$ system at 2400K: Iso-surfaces of electron density obtained by DFT (a) and ML model for original (unpruned) (b), 90% CCS (c), and 99.5% CCS (d). The blue, purple, and green spheres represent the Si, Ge, and Sn atoms.
Figure 25: $\text{Si}_{50}\text{Ge}_{12.5}\text{Sn}_{37.5}$ system at 2400K: Iso-surfaces of errors in electron density obtained by ML model for original (a), 90% CCS (b), and 99.5% CCS (c) respectively.
Figure 26: $\text{Cr}_{25}\text{Fe}_{25}\text{Co}_{25}\text{Ni}_{25}$ system at 5000K: Iso-surfaces of electron density obtained by DFT (a) and ML model for original (unpruned) (b), 90% CCS (c), and 99.5% CCS (d). The green, orange, blue, and gray spheres represent the Cr, Fe, Co, and Ni atoms.
Figure 27: $\text{Cr}_{25}\text{Fe}_{25}\text{Co}_{25}\text{Ni}_{25}$ system at 5000K: Iso-surfaces of errors in electron density obtained by ML model for original (a), 90% CCS (b), and 99.5% CCS (c) respectively.